if 1+sin a =9 cos a then 5(1-cos a)=
  1. sin a 
  2. 2 sin a
  3. 4 isn a
  4. tan a

Grade:11

1 Answers

Vikas TU
7 years ago
 1+sin a =9 cos a
Divide by cosa to the whole eqn.,
seca + tana = 9
seca = 9 – tana
sq. both sides,
sec^2a = 81 + tan^2a  – 18tana
1 + tan^2a = 81 – 18tana + tan^2a
tan^2a got cancel.
80 = 18tana
tana = 80/18
tana = 40/9
Now find cosa and cross check the options now.